# Full-Duplex-Bench: A Benchmark to Evaluate Full-duplex Spoken Dialogue Models on Turn-taking Capabilities

### v1

@@ Abstract
Spoken dialogue modeling introduces unique challenges beyond text-based language modeling, demanding robust turn-taking, backchanneling, and real-time interaction. Although most Spoken Dialogue Models (SDMs) rely on half-duplex processing (handling speech one turn at a time), emerging full-duplex SDMs can listen and speak simultaneously, enabling more natural and engaging conversations. However, current evaluations of such models remain limited, often focusing on turn-based metrics or high-level corpus analyses (e.g., turn gaps, pauses). To address this gap, we present Full-Duplex-Bench [^1] Data, codebase, and demo samples will be released here https://github.com/DanielLin94144/Full-Duplex-Bench., a new benchmark that systematically evaluates key conversational behaviors: pause handling, backchanneling, turn-taking, and interruption management. Our framework uses automatic metrics for consistent and reproducible assessments of SDMs’ interactive performance. By offering an open and standardized evaluation benchmark, we aim to advance spoken dialogue modeling and encourage the development of more interactive and natural dialogue systems.
Full-Duplex-Bench: A Benchmark to Evaluate Full-duplex Spoken Dialogue Models on Turn-taking Capabilities
Guan-Ting Lin1  Jiachen Lian2thanks: Equal contribution, listed alphabetically.  Tingle Li2∗ Qirui Wang3∗ Gopala Anumanchipalli2  Alexander H. Liu4  Hung-yi Lee1 1Graduate Institute of Communication Engineering, National Taiwan University 2UC Berkeley  3University of Washington  4MIT CSAIL
1 Introduction
Natural spoken dialogue is characterized by complex dynamics (Schegloff, 2000; Gravano and Hirschberg, 2011). Different from text-based language modeling, spoken dialogue involves problems beyond language understanding for communication, including turn-taking (Duncan, 1972; Gravano and Hirschberg, 2011; Raux and Eskenazi, 2012), and backchanneling (Schegloff, 1982). These fundamental aspects facilitate mutual understanding, engagement, and social connection. Over the years, Spoken Dialogue Models (SDMs) have become a major research focus, particularly with the rise of voice assistants such as Apple Siri, Google Assistant, and Amazon Alexa.
Figure 1: Overview pipeline of Full-Duplex-Bench. We feed user audio streams to an open-source full-duplex SDM, which produces time-synchronous output. We then perform post-processing to align both streams at the transcript level, enabling automatic evaluation along multiple dimensions.
@@ 1 Introduction
Although half-duplex SDMs have been the dominant approach, they often sound less natural than human speakers, who seamlessly listen and speak at the same time. Recent progress in models like GPT-4o voice mode has sparked an interest in full-duplex capabilities for more human-like dialogues. Consequently, many new SDMs have emerged (Veluri et al., 2024; Défossez et al., 2024; Wang et al., 2024a; Fu et al., 2024; Wang et al., 2024c; Chen et al., 2025; Ji et al., 2024), with different architectures to enable real-time communication. As an increasing number of these systems are proposed, fair and open evaluation is critical to guide future advances in spoken dialogue research.
@@
Evaluating SDMs is challenging because human conversation requires understanding a wide range of information and interpreting complex interactions within speech signals. Previous benchmarks have largely focused on: Content-based evaluation, measuring performance on tasks such as spoken question answering (Hassid et al., 2024; Nachmani et al., 2024; Shih et al., 2024; Lin et al., 2024c, 2022); Instruction-following evaluation, assessing a system’s ability to follow directives (Huang et al., 2024b, a); and Paralinguistic-based evaluation, examining factors like emotion or speaking style (Cheng et al., 2025; Ao et al., 2024; Yang et al., 2024; Lin et al., 2024a, b; Lin and Lee, 2024). However, these benchmarks mostly assume turn-based interactions, leaving the real-time aspects of full-duplex models under-explored.
@@
Some recent studies attempt to evaluate interaction timing in full-duplex dialogue. dGSLM (Nguyen et al., 2023) examines voice activity patterns by comparing model outputs to ground truth gaps, interpausal units, and pauses. However, the corpus-level statistics are difficult to interpret, and the ground truth is based on a specific dialogue dataset. Recently, Talking-Turns (Arora et al., 2025) uses a specialized judge model to evaluate alignment with conversational behaviors, but the model is trained on a specific dialogue dataset, which may not generalize to different dialogues. While these works offer valuable insights, they emphasize the need for a more unified and reproducible approach to testing full-duplex SDMs in real-life scenarios.
@@
In this paper, we address this gap by introducing the first benchmark designed to systematically evaluate key turn-taking behaviors in full-duplex SDMs (Figure 1). Our framework focuses on four crucial aspects – pause handling, backchanneling, turn-taking, and user interruption management, and uses automatic metrics for reliable comparisons. By evaluating our benchmark to SDMs, we can better understand how these systems handle real-time interactive dynamics. We will publicly release the benchmark and the associated codebase to encourage the further development of more engaging and conversational SDMs.
2 Related Works
2.1 Full-duplex Spoken Dialogue Models
Full-duplex SDMs aim to mimic the natural flow of human conversation by allowing simultaneous listening and speaking. Recent approaches can be divided into two main categories:
End-to-End Two-Channel Models
dGSLM (Nguyen et al., 2023) applies a Siamese network with cross-attention to handle two-channel spoken dialogue. Moshi (Défossez et al., 2024) uses parallel processing streams for speaker and listener speech. SyncLLM (Veluri et al., 2024) introduces a time-synchronous mechanism for autoregressive decoding across both channels. OmniFlatten (Zhang et al., 2024a) employs progressive post-training to jointly process flattened sequences of speech and text tokens, while SALMONN-omni (Yu et al., 2024) and MinMo (Chen et al., 2025) augment large language models with additional state tokens to refine speech-based intelligence. Parrot (Wang et al., 2025) uses next-token-pair prediction for two-channel modeling. RTTL-DG (Mai and Carson-Berndsen, 2025) leverages two-channel input with a dialogue manager to predict the speaking or listening state to determine when to generate the response.
Cascaded Models
Wang et al. (2024b) introduce control, speak, and listen tokens on top of a large language model with perception and motor modules, although their design still relies on speech-to-text and text-to-speech components. MiniCPM-Duplex (Zhang et al., 2024b) and MiniCPM-Duo (Xu et al., 2024) incorporate time-sliced chunks of text tokens, adapting time-synchronous methods for language modeling. VITA (Fu et al., 2024) and Freeze-Omni (Wang et al., 2024c) accept raw speech input directly, improving inference speed but still requiring text in the loop. These varied strategies highlight a clear need for standardized and open benchmarks that can evaluate turn-taking and real-time interaction across emerging full-duplex SDMs.
2.2 Evaluation Benchmarks
Benchmarks offer a common framework for comparing models and understanding their performance. In the speech community, several benchmarks have been proposed to evaluate specific aspects of speech models. The SUPERB series (wen Yang et al., 2021; Lin et al., 2023; Tsai et al., 2022) evaluates self-supervised speech representations across a variety of tasks, while ML-SUPERB (Shi et al., 2023), IndicSUPERB (Javed et al., 2023), and LeBenchmark (Evain et al., 2021) extend these evaluations to multilingual and non-English models.
@@ 2.2 Evaluation Benchmarks
Recent advances in universal spoken language models have resulted in benchmarks that evaluate both understanding and generation. These can be grouped as follows: Multi-task Instruction Following: Dynamic-SUPERB (Huang et al., 2024b, a), Voicebench (Chen et al., 2024), and AIR-Bench (Yang et al., 2024) test a model’s ability to perform a range of tasks based on specific instructions. Semantic Understanding: For low-level language skills such as syntax and grammar, benchmarks like Zerospeech (Nguyen et al., 2020) and datasets such as Spoken-StoryCloze (Mostafazadeh et al., 2017) have been introduced. Higher-level comprehension, including spoken question answering, is measured by benchmarks like Llama-Question (Nachmani et al., 2024) and Spoken WebQuestion. Paralinguistic Perception: VoxDialogue (Cheng et al., 2025), SD-Eval (Ao et al., 2024), and StyleTalk (Lin et al., 2024a) evaluate the ability to produce natural responses by considering features such as emotion, gender, and accent.
Despite these developments, no existing benchmark explicitly evaluates the turn-taking abilities of SDMs, a critical aspect of natural spoken interaction. To address this gap, we introduce a new benchmark that measures SDMs’ turn-taking behavior across multiple dimensions. This work propose a new benchmark for evaluating conversational fluidity, paving the way for the development of more interactive and natural dialogue systems.
3 Full-Duplex-Bench Framework
3.1 Overview
Our methodology utilizes unified speech input to simulate real-time user interactions with SDMs, creating controlled scenarios for each evaluation dimension. We prepare specific test samples for each aspect, collect the model’s speech responses, and apply CrispyWhisper (Zusag et al., 2024) to generate time-aligned transcriptions. Specialized metrics are then employed to assess performance in each dimension.
3.2 Prerequisite
@@
Backchannel
In natural conversations, listeners often provide short verbal or nonverbal signals, known as backchannels, to indicate active engagement and understanding. These backchannels include utterances such as “mm-hmm,” and “uh-huh", which typically occur while another person is speaking. Effective backchanneling enhances conversational fluidity by signaling attentiveness without disrupting the speaker. Generally, backchanneling refers to short utterances produced by the listener while the speaker is talking.
In this work, we classify a speech segment as backchanneling if it meets the following criteria: (1) it has a short duration of less than 1 second; and (2) it contains fewer than two words. This ensures that speech is delivered at a reasonable pace and that brief utterances do not interrupt the current speaker’s turn. Although the concept of backchanneling differs across the literature (Ward and Tsukahara, 2000; Reece et al., 2023), developing a comprehensive detector remains a goal for future work.
Takeover (TO)
A takeover occurs when the model effectively assumes control of the conversation, dominating the turn and granting minimal opportunity for the user to speak. In this work, takeover is treated as a binary variable. If the model merely responds with silence or a backchannel, no takeover is deemed to have occurred. Conversely, any other non-silent speech that is not a backchannel indicates the model’s attempt to take over. Formally,
@@
Takeover Rate (TOR)
To quantify how often takeovers occur, we define the TOR as the average value of the binary TO variable across the dataset: $\text{TOR}=\frac{1}{N}\sum_{i=1}^{N}\text{TO}_{i}$, where $N$ is the total number of samples, and $\text{TO}_{i}$ is the binary takeover variable for sample $i$.
3.3 Evaluation Dimensions
@@ 3.3.1 Pause Handling
Research Question
Can the model recognize when the other speaker is still holding the turn and understand that it should not take over?
Metric
The ideal model behavior is to avoid taking over the turn while the user is speaking. To evaluate this, we use the Takeover Rate (TOR). A lower TOR signifies better pause management, indicating that the model effectively waits for the user’s turn to end. In contrast, a higher TOR suggests that the model is more likely to take over the conversation before the user has yielded the turn.
3.3.2 Backchanneling
Drawing on the definition in Section 3.2, we assess whether the model, when interacting with a dominant speaker, actively listens and provides backchannels at suitable moments to facilitate dialogue engagement. A model exhibiting human-like backchanneling behavior should respond at the right times and with an appropriate frequency.
Research Question
Can the model determine when to offer backchannels in a human-like manner without interrupting the speaker?
Metric
To measure how well models predict the timing of human backchannel cues (e.g., “uh-huh,” “mm-hmm”) during conversations, we use three metrics:
@@ 3.3.2 Backchanneling
  • Backchannel Frequency (Freq): Each backchannel event is counted and normalized by duration (events per second). When the model does not take over the turn (TOR = 0), a higher frequency suggests more effective backchanneling.
  • Jensen-Shannon Divergence (JSD): This captures the difference between the model’s predicted timing of backchannels and actual human timing. The model’s predictions form a probability distribution $P$, indicating the likelihood of a backchannel within each discrete time window. Similarly, the human ground truth forms a distribution $Q$ over the same set of time windows. We compute $\text{JSD}(P||Q)=\frac{1}{2}D_{\text{KL}}(P||M)+\frac{1}{2}D_{\text{KL}}(Q||M)$, where $M=\frac{P+Q}{2}$ is the average distribution, and $D_{\text{KL}}$ is the Kullback–Leibler divergence. Expanded, $\text{JSD}(P||Q)=\frac{1}{2}\sum_{i}P(i)\log\frac{P(i)}{M(i)}+\frac{1}{2}\sum_{i}Q(i)\log\frac{Q(i)}{M(i)}$. JSD ranges from 0 (perfect agreement) to 1 (complete divergence). A lower score indicates closer alignment with human backchannel timing. We only calculate this metric when the model does not take over the turn, where each backchannel event is counted as one-hot and normalized into a probability distribution. If the model stays silent throughout, we assume a uniform probability distribution, treating it as a random baseline without backchannel knowledge.
3.3.3 Smooth Turn Taking
@@
Research Question
Can the model detect the end of a speaker’s turn and respond promptly without long pauses?
Metric
We measure the averaged response latency, the time (in seconds) between the end of the user’s speech and the start of the model’s response. Lower latency values indicate smoother turn-taking. In cases where the model fails to respond, we record the TOR. The latency is calculated only when TO equals 1. This avoids averaging with non-takeover periods, which would introduce significant variance due to periods of silence.
3.3.4 User Interruption
@@
Research Question
Can the model detect and adapt to user interruptions while maintaining a coherent and timely response?
Metric
The evaluation consists of three measures:
@@
  • GPT-4o Score: A large language model evaluates the quality of the system’s response (coherence, relevance, and adaptability) on a scale from 0 to 5. Higher scores indicate better performance. Detailed instructions are provided in Appendix Table 5.
@@
Model Date E2E #ch Interrupt BC S2S Release
Transparent Models
dGSLM (Nguyen et al., 2023) 2022/3 2
FSM (Wang et al., 2024b) 2024/5 1 - -
MiniCPM-Duplex (Zhang et al., 2024b) 2024/6 1 - -
VITA (Fu et al., 2024) 2024/8 1 -
SyncLLM Veluri et al. (2024) 2024/9 2
Parrot Wang et al. (2025) 2024/9 2 -
MiniCPM-Duo (Xu et al., 2024) 2024/9 1 - -
Moshi (Défossez et al., 2024) 2024/10 2
SALMONN-omni (Yu et al., 2024) 2024/11 1 -
MinMo (Chen et al., 2025) 2025/1 1
OmniFlatten (Zhang et al., 2024a) 2025/1 2 -
RTTL-DG (Mai and Carson-Berndsen, 2025) 2025/1 2
Freeze-Omni Wang et al. (2024c) 2024/11 1 -
Closed-source Commercial Models
GPT-4o Voice Mode 2024/5 - - -
Gemini Live Voice Chat 2024/8 - - -
DouBao 2025/1 - - -
Table 1: Overview of full-duplex speech language models. “—” indicates properties that are not publicly available or unspecified. “E2E” denotes end-to-end speech modeling (without relying on text), “#ch” indicates the number of speech input channels, “Interrupt” refers to interruption handling, “BC” represents backchanneling capability, and “S2S Release” shows whether the complete speech-to-speech pipeline is publicly released (some models only release partial checkpoints without this functionality).
3.4 Data Curation

### v2

@@ Abstract
Spoken dialogue modeling poses challenges beyond text-based language modeling, requiring real-time interaction, turn-taking, and backchanneling. While most Spoken Dialogue Models (SDMs) operate in half-duplex mode—processing one turn at a time—emerging full-duplex SDMs can listen and speak simultaneously, enabling more natural conversations. However, current evaluations remain limited, focusing mainly on turn-based metrics or coarse corpus-level analyses. To address this, we introduce Full-Duplex-Bench [^1] Data, codebase, and demo samples will be released here https://github.com/DanielLin94144/Full-Duplex-Bench., a benchmark that systematically evaluates key interactive behaviors: pause handling, backchanneling, turn-taking, and interruption management. Our framework uses automatic metrics for consistent, reproducible assessment and provides a fair, fast evaluation setup. By releasing our benchmark and code, we aim to advance spoken dialogue modeling and foster the development of more natural and engaging SDMs.
Index Terms:
Full-Duplex Dialogue, Benchmark, Spoken Language Models, Turn-Taking
I Introduction
Natural spoken dialogue is characterized by complex dynamics [1, 2]. Different from text-based language modeling, spoken dialogue involves problems beyond language understanding for communication, such as turn-taking [3, 2, 4], and backchanneling [5]. These fundamental aspects facilitate mutual understanding, engagement, and social connection. Over the years, Spoken Dialogue Models (SDMs) have become a major research focus, particularly with the rise of voice assistants.
Figure 1: Overview pipeline of Full-Duplex-Bench. We feed user audio streams to a full-duplex SDM, which produces time-synchronous output. We then perform post-processing to align both streams at the transcript level, enabling automatic evaluation along multiple dimensions.
@@ I Introduction
Although half-duplex SDMs have been the dominant approach, they often sound less natural than human speakers, who seamlessly listen and speak at the same time. Recent progress in models like GPT-4o voice mode has sparked an interest in full-duplex capabilities for more human-like dialogues. Consequently, many new SDMs have emerged [6, 7, 8, 9, 10, 11, 12], with different architectures to enable real-time communication. As an increasing number of these systems are proposed, fair and open evaluation is critical to guide future advances in spoken dialogue research.
@@
Evaluating SDMs is challenging because human conversation requires understanding a wide range of information and interpreting complex interactions within speech signals. Previous benchmarks have largely focused on: Content-based evaluation, measuring performance on tasks such as spoken question answering [13, 14, 15, 16, 17]; Instruction-following evaluation, assessing a system’s ability to follow directives [18, 19]; and Paralinguistic evaluation, examining factors like emotion or speaking style [20, 21, 22, 23, 24, 25]. However, these benchmarks mostly assume turn-based interactions, leaving the real-time aspects of full-duplex models under-explored.
TABLE I: Overview of full-duplex speech language models. “—” indicates properties that are not publicly available or unspecified. “E2E” denotes end-to-end speech modeling (without relying on text), “#ch” indicates the number of speech input channels, “Interrupt” refers to interruption handling, “BC” represents backchanneling capability, and “S2S Release” shows whether the complete speech-to-speech pipeline is publicly released (some models only release partial checkpoints without this functionality).
Model Date E2E #ch Interrupt BC S2S Release
Transparent Models
dGSLM [26] 2022/3 2
FSM [27] 2024/5 1 - -
MiniCPM-Duplex [28] 2024/6 1 - -
VITA [9] 2024/8 1 -
SyncLLM [6] 2024/9 2
Parrot [29] 2024/9 2 -
MiniCPM-Duo [30] 2024/9 1 - -
Moshi [7] 2024/10 2
SALMONN-omni [31] 2024/11 1 -
MinMo [11] 2025/1 1
OmniFlatten [32] 2025/1 2 -
RTTL-DG [33] 2025/1 2
Freeze-Omni [10] 2024/11 1 -
Closed-source Commercial Models
GPT-4o Voice Mode 2024/5 - - -
Gemini Live 2024/8 - - -
DouBao 2025/1 - - -
Nova Sonic [34] 2025/4 - - -
@@
Some recent studies attempt to evaluate interaction timing in full-duplex dialogue. dGSLM [26] examines voice activity patterns by comparing model outputs to ground truth gaps, interpausal units, and pauses. However, the corpus-level statistics are difficult to interpret, and the ground truth is based on a specific dialogue dataset. Recently, Talking-Turns [35] trains a specialized judge model to evaluate alignment with conversational behaviors. However, this model is trained on a specific dialogue dataset, which may limit its generalizability to other dialogue types. Moreover, the evaluation relies on user studies, making the results difficult to reproduce due to variations in participants. While these works provide valuable insights, they highlight the need for a more unified and reproducible approach to evaluating full-duplex SDMs in real-world scenarios.
@@
In this paper, we address the need by introducing Full-Duplex-Bench, the first benchmark designed to systematically evaluate key turn-taking behaviors in full-duplex spoken dialogue models (SDMs), as illustrated in Figure 1. Our framework targets four critical aspects of real-time interaction: pause handling, backchanneling, turn-taking, and user interruption management, using automatic metrics to enable consistent and reliable comparisons. This benchmark represents an important initial step toward evaluating the full-duplex turn-taking capabilities of SDMs. Moreover, the benchmark can be executed efficiently and automatically, enabling rapid evaluation across models. By applying our benchmark to existing systems, we gain deeper insights into how they manage interactive dynamics in real time. We will publicly release the benchmark and associated codebase to encourage further research and the development of more engaging and conversational SDMs.
II Related Works
II-A Full-duplex Spoken Dialogue Models
Recent advances in SDMs have increasingly focused on full-duplex capabilities. We categorize existing full-duplex SDMs into two main groups: Transparent Models, which offer detailed architecture and implementation, and Closed-Source Commercial Systems, which are accessible only through demos. Table I summarizes the key characteristics of these models.
Transparent Models: These models are publicly available with open-source implementations or detailed technical descriptions, enabling reproducibility and deeper analysis. They fall into two subcategories:
Cascaded Models. These systems follow a modular pipeline structure, typically integrating ASR, LLM, and TTS components. Wang et al. [27] introduce control, speak, and listen tokens in an LLM framework with perceptual inputs. MiniCPM-Duplex [28] and MiniCPM-Duo [30] apply time-sliced token windows for synchronous dialogue modeling. VITA [9] and Freeze-Omni [10] operate on raw speech inputs, improving latency while maintaining internal reliance on text. While cascaded approaches offer flexibility and stronger semantic modeling, they often suffer from cross-module latency and the loss of nuanced speech features, driving growing interest in end-to-end alternatives.
End-to-End Models. These models jointly model speaker and listener audio streams. dGSLM [26] employs a Siamese network with cross-attention for two-channel dialogue. Moshi [7] uses parallel processing streams to handle real-time user speech and supports overlaps and interruptions. SyncLLM [6] introduces time-synchronous audio chunks modeling for two-channel speech. OmniFlatten [32] flattens and processes speech and text tokens jointly. SALMONN-omni [31] and MinMo [11] inject state tokens to improve turn-taking modeling. Parrot [29] predicts next-token pairs across streams, while RTTL-DG [33] uses a dialogue manager to control generation timing. By operating directly on audio data, end-to-end approaches have the potential to capture a wider range of speech features—including paralinguistic and non-verbal cues—as well as dialogue behaviors such as backchanneling.
Closed-Source Commercial Systems: Several commercial models have demonstrated full-duplex capabilities through public demos. These systems—such as GPT-4o Voice Mode [^2] https://openai.com/index/hello-gpt-4o/, Gemini Live [^3] https://blog.google/products/gemini/made-by-google-gemini-ai-updates/, Doubao [^4] https://team.doubao.com/en/special/realtime_voice, and Nova Sonic [34]—are not open-sourced and lack publicly available architectural details. Nonetheless, we include them in Table I. Their emergence underscores the need for open and standardized benchmarks to evaluate interaction quality across both transparent research models and proprietary commercial systems.
II-B Evaluation Benchmarks
@@
Benchmarks offer a common framework for comparing models and understanding their performance. In the speech community, several benchmarks have been proposed to evaluate specific aspects of speech models. The SUPERB series [36, 37, 38] evaluates self-supervised speech representations across a variety of tasks. Recent advances in universal spoken language models have resulted in benchmarks that evaluate both understanding and generation. These can be grouped as follows: Multi-task Instruction Following: Dynamic-SUPERB [18, 19], Voicebench [39], and AIR-Bench [22] test a model’s ability to perform a range of tasks based on specific instructions. Semantic Understanding: For low-level language skills such as syntax and grammar, benchmarks like Zerospeech [40] and datasets such as Spoken-StoryCloze [41] have been introduced. Higher-level comprehension, including spoken question answering, is measured by benchmarks like Llama-Question [14] and Spoken WebQuestion. Paralinguistic Perception: VoxDialogue [20], SD-Eval [21], and StyleTalk [23] evaluate the ability to produce natural responses by considering features such as emotion, gender, and accent. Talking-Turns [35] trains a judge model on a specific dialogue dataset to assess conversational alignment, but its generalizability is limited. Its reliance on user studies also hinders reproducibility. These limitations highlight the need for a unified, reproducible framework for evaluating full-duplex SDMs. To this end, we propose a benchmark with an automatic evaluation pipeline and behavior-specific metrics for assessing conversational performance.
III Full-Duplex-Bench Framework
@@
Backchannel: In natural conversations, listeners often provide short verbal or nonverbal signals, known as backchannels, to indicate active engagement and understanding. These backchannels include utterances such as “mm-hmm,” and “uh-huh", which typically occur while another person is speaking. Effective backchanneling enhances conversational fluidity by signaling attentiveness without disrupting the speaker. Generally, backchanneling refers to short utterances produced by the listener while the speaker is talking.
In this work, we classify a speech segment as backchanneling if it meets the following criteria: (1) it has a short duration of less than 1 second; and (2) it contains fewer than two words. This ensures that speech is delivered at a reasonable pace and that brief utterances do not interrupt the current speaker’s turn. The concept of backchanneling differs across the literature [42, 43], so developing a comprehensive detector is beyond the scope of this paper and is left for future work.
Takeover (TO): A takeover occurs when the model effectively assumes control of the conversation, dominating the turn and granting minimal opportunity for the user to speak. In this work, takeover is treated as a binary variable. If the model merely responds with silence or a backchannel, no takeover is deemed to have occurred. Conversely, any other non-silent speech that is not a backchannel indicates the model’s attempt to take over. Formally,
@@
Takeover Rate (TOR): To quantify how often takeovers occur, we define the TOR as the average value of the binary TO variable across the dataset: $\text{TOR}=\frac{1}{N}\sum_{i=1}^{N}\text{TO}_{i}$, where $N$ is the total number of dialogue turns, and $\text{TO}_{i}$ is the binary takeover variable for sample $i$.
III-A Overview
As illustrated in Figure 1, our framework employs a unified speech input (denoted as “input.wav") to simulate real-time user interactions with SDMs, enabling controlled evaluations across various dimensions. For each aspect, we design targeted test samples, collect the model’s speech responses (“output.wav"), and use an ASR model (Nvidia parakeet-tdt-0.6b-v2555https://huggingface.co/nvidia/parakeet-tdt-0.6b-v2 to produce word-level, time-aligned transcriptions (“output.json"). Dedicated metrics are then applied to assess performance in each dimension.
III-B Evaluation Dimensions
@@
Research Question: Can the model recognize when the other speaker is still holding the turn and understand that it should not take over?
Metric: The ideal model behavior is to avoid taking over the turn while the user is speaking. To evaluate this, we use the Takeover Rate (TOR). A lower TOR signifies better pause management, indicating that the model effectively waits for the user’s turn to end. In contrast, a higher TOR suggests that the model is more likely to take over the conversation before the user has yielded the turn.
2) Backchanneling
Drawing on the definition in Section III, we assess whether the model, when interacting with a dominant speaker, actively listens and provides backchannels at suitable moments to facilitate dialogue engagement. A model exhibiting human-like backchanneling behavior should respond at the right times and with an appropriate frequency.
Research Question: Can the model determine when to offer backchannels in a human-like manner without interrupting the speaker?
Metric: To measure how well the models generate backchannel cues, we use three metrics:
@@
  • Backchannel Frequency (Freq): Each backchannel event is counted and normalized by duration (events per second). When the model does not take over the turn (TOR = 0), a higher backchanneling frequency indicates that the model responds backchannel more often, but this does not necessarily imply better or more natural behavior, as it also depends on timing and context.
  • Jensen-Shannon Divergence (JSD): This captures the difference between the model’s predicted timing of backchannels and actual human timing. The model outputs a probability distribution $P$, where $P(i)$ denotes the likelihood of a backchannel occurring in time window $i$. The ground truth distribution $Q$ is derived from human-annotated backchannel timings (details in III-C), aligned to the same set of time windows. To measure the similarity between $P$ and $Q$, we compute the Jensen–Shannon Divergence (JSD) as:
    $\text{JSD}(P||Q)=\frac{1}{2}\sum_{i}P(i)\log\frac{P(i)}{M(i)}+\frac{1}{2}\sum_{i}Q(i)\log\frac{Q(i)}{M(i)},$
    where $M(i)=\frac{1}{2}(P(i)+Q(i))$, and $i$ indexes the discrete time windows. JSD ranges from 0 (perfect alignment) to 1 (complete divergence), providing a symmetric and bounded measure of similarity between model predictions and human backchannel behavior. We only calculate this metric when the model does not take over the turn, where each backchannel event is counted as one-hot and normalized into a probability distribution. If the model stays silent throughout, we assume a uniform probability distribution, treating it as a random baseline without backchannel knowledge.
3) Smooth Turn Taking
@@
Research Question: Can the model detect the end of a speaker’s turn and respond promptly without long pauses?
Metric: We measure the averaged response latency, the time (in seconds) between the end of the user’s speech and the start of the model’s response. Lower latency values indicate smoother turn-taking. In cases where the model fails to respond, we record the TOR. The latency is calculated only when TO equals 1. This avoids averaging with non-takeover periods, which would introduce significant variance due to periods of silence.
4) User Interruption
@@
Research Question: Can the model detect and adapt to user interruptions while maintaining a coherent and timely response?
Metric: The evaluation consists of three measures:
@@
  • GPT-4o Score: A large language model evaluates the quality of the system’s response (coherence, relevance, and adaptability) on a scale from 0 to 5. Higher scores indicate better performance.
@@
TABLE II: The number of samples for each dataset and task in Full-Duplex-Bench.
Dataset Task # of Samples
Candor Pause Handling 216
Candor Smooth Turn-Taking 119
ICC Backchannel 55
Synthetic User Interruption 200
Synthetic Pause Handling 137
III-C Data Curation